P=100-2Q where Q is total quantity demanded for both firms 1 and 2, respectively. The firms 1 marginal cost is given by MC1(Q1)=2Q1. The firms 2 marginal cost is given by MC2(Q2)=4Q2. Based on this information firm 1 and firm 2's reaction functions are?
Answer #1

P- 100 - 20,-2a2 From 1 TR, = lood, -29, -29,02 MR, = 100-49,- 20₂ 2 Me = Mira MC 100-49, -222 =2 98 - 292 = 40 9,= 24.5 - 1